OpenStudy (anonymous):

What is the solution set for |–2x + 16| – 9 = –3 is it 5 and 11 or 5 and -11

OpenStudy (anonymous):

if |-2x+16| > o the equation stays the same. but if it is <0 : 2x-16 -9=-3

OpenStudy (anonymous):

|–2x + 16| – 9 = –3 |–2x + 16| = –3 +9 –2x + 16 = 6 –2x + 16 = -6 Solve them...

OpenStudy (anonymous):

ok c: